# Visual Structures Helps Visual Reasoning: Addressing the Binding Problem in VLMs

### v2

@@ 4.2 Visual Search
Dataset Objects Simple Ours Simple Ours Simple Ours Simple Ours
2D 20 0.71 0.91 0.55 0.82 0.00 0.06 0.31 0.43
30 0.50 0.80 0.54 0.70 0.00 0.00 0.28 0.37
40 0.25 0.64 0.18 0.59 0.00 0.00 0.14 0.23
50 0.46 0.55 0.10 0.55 0.00 0.00 0.47 0.56
Avg 0.48 0.73 0.34 0.66 0.00 0.02 0.30 0.40
3D 20 1.00 1.00 1.00 1.00 0.82 0.80 0.33 0.44
30 0.96 0.96 0.88 0.91 0.43 0.60 0.15 0.18
40 0.89 0.94 0.72 0.80 0.59 0.58 0.00 0.15
50 0.81 0.84 0.59 0.75 0.39 0.38 0.00 0.04
Avg 0.91 0.93 0.80 0.86 0.56 0.59 0.12 0.20
An instance of the visual search problem is presented in Figure 2-(a); to evaluate our method, we use the synthetic dataset described earlier to generate 2D and 3D scenes containing between 20 and 50 objects of varying complexities. We assessed performance by measuring accuracy in correctly identifying the target’s presence or absence, distinguishing between “visible” searches (the object is explicitly in the scene) and “invisible” searches (the object is absent). This distinction is important because simpler visual-language models often default to predicting “false,” inflating their invisible-object accuracy while failing to detect visible targets. After computing the accuracy for visible and invisible subtasks, we combined them via the harmonic mean to produce a final score.
Harmonic $\text{Harmonic Mean}=\frac{2\cdot\text{Visible}\cdot\text{Invisible}}{\text{Visible}+\text{Invisible}}$ (1)
Table 1 compares our approach against baselines, showing our method delivers consistent and substantial gains over the simple baseline across 2D and 3D scenes and for all evaluated VLMs. In 2D settings, the average harmonic mean for GPT-4o climbs from 0.48 to 0.73, for Claude3.5-sonnet from 0.34 to 0.66, for LLaMa4 from 0.00 to 0.02, and for Qwen2.5-VL from 0.30 to 0.40, with the largest relative improvements occurring as scene complexity increases (e.g., at 40 objects GPT-4o rises from 0.25 to 0.64, and Claude3.5-sonnet rises from 0.18 to 0.59). In 3D scenes our method still yields gains to 0.93, and weaker models benefit notably (Claude3.5-sonnet: 0.80 → 0.86; LLaMa4: 0.56 → 0.59; Qwen2.5-VL: 0.12 → 0.20). These results confirm that our strategy not only corrects the “always-false” bias of simpler VLMs in invisible-object tasks but also markedly enhances their ability to detect visible targets, achieving robust performance regardless of object count, visibility, or scene dimensionality.
4.3 Counting
Counting involves determining the number of specific objects within a scene. Although it may appear simple, accurate enumeration—especially when multiple object types or attribute variations are involved—relies heavily on effective feature binding. Each object must be individuated, its defining features correctly associated, and then counted as a distinct instance. Failures in binding can lead to omissions, double-counting, or misclassification. Similar to human rapid estimation under time constraints, VLMs show capacity limits in counting. Performance often worsens as object numbers grow, or when high object similarity (low feature heterogeneity) increases the interference and binding errors.
Table 2: Performance comparison on the counting task using models GPT-4o, Claude3.5-sonnet, LLaMa4, and Qwen2.5-VL, evaluated with accuracy (%) . The comparison is conducted across 2D and 3D datasets with varying numbers of objects in the scene, as well as a natural image dataset. Results are shown for base models and our method.
GPT-4o Claude-sonnet LLaMa4 Qwen2.5-VL
Dataset Objects Simple Ours Simple Ours Simple Ours Simple Ours
@@ 4.3 Counting
An instance of this task is demonstrated in Figure 2-(c); To evaluate our method, we generated 2D and 3D images, using the synthetic data set described above, each containing between 10 and 20 instances of the target object to increase complexity. We measure performance using counting accuracy; the results are summarized in Table 2. As presented, our counting-specific prompting yields substantial gains on synthetic benchmarks: in 2D scenes, GPT-4o jumps from 12.0% to 38.8%, Claude-sonnet from 8.5% to 10.67%, LLaMa4 from 6.3% to 19.1%, and Qwen2.5-VL surges from 5.8% to 40.8%; in 3D scenes, GPT-4o climbs from 15.0% to 31.0%, Claude-sonnet from 17.3% to 22.2%, LLaMa4 from 14.0% to 36.4%, and Qwen2.5-VL from 8.5% to 26.7%. On natural images, improvements are more modest—GPT-4o from 29.8% to 35.7%, Claude-sonnet from 2.09% to 6.4%, LLaMa4 from 29.2% to 31.7%—and Qwen2.5-VL sees a slight drop (18.9% to 17.3%). These results underscore the effectiveness of task-aware prompting in controlled settings while highlighting the remaining challenge of generalizing to real-world imagery.
@@ 4.4 Scene Description
The scene description task requires models to generate accurate textual narratives that describe the objects, their attributes, and their relationships within an image. This task challenges a VLM’s ability to solve the binding problem. The model must not only detect features (e.g., “red”, “cube”, “blue”, “sphere”) but also correctly associate them with the corresponding objects in its textual output (e.g., “a red cube and a blue sphere”, not “a blue cube and a red sphere”). Errors, such as attribute swapping or incorrectly pairing features with objects, often arise in complex scenes. This typically occurs when multiple objects share some features but differ in others, which inherently challenges precise feature-object binding.
Table 3: Comparison of scene description performance (lower is better; measured as average distance) for our method versus base models and CoT prompting across GPT-4o and Claude3.5-sonnet on 2D and 3D datasets with varying numbers of objects.
@@
Dataset Objects Simple Ours Simple Ours Simple Ours Simple Ours
@@
Figure 2-(b) shows an instance of the scene description task; we evaluate a synthetically generated dataset of 2D and 3D scenes, each containing between 10 and 20 objects to vary scene complexity. Model performance is quantified by the edit distance between the generated descriptions and the ground-truth annotations; we show more detail about this in Appendix C. This yields three outcome categories: (1) an exact match, (2) a single-attribute mismatch (either shape or color), and (3) a double-attribute mismatch (both shape and color). The full results are presented in Table 3. The greatest average gains (-0.81 in 2D and -1.06 in 3D) are driven precisely where the task’s difficulty peaks, showing that the technique not only endures extra clutter but thrives on it. This confirms that our method is the most reliable option for reasoning with a high object count.
@@ 4.5 Spatial Relationship
In this task, the model is expected to identify and verify the relative positions of objects in a scene (e.g., “Is the red cube to the left of the blue sphere?”). This involves not only correctly binding intrinsic features (e.g., color, shape) to objects but also accurately associating each object with its spatial location, followed by a relational comparison. Failure to bind features properly can lead to misidentification of the objects being queried. Furthermore, if the VLM cannot accurately associate each identified object with its precise location, its capacity to reason about their relative positions will be compromised. VLMs often struggle with spatial reasoning due to their inability to maintain distinct, spatially grounded representations of multiple objects simultaneously [8].
Table 4: Accuracy (%) comparison of spatial relationship predictions across GPT-4o, Claude3.5-sonnet, LLaMa4, and Qwen2.5-VL on 2D, and natural image datasets using base models, CoT prompting, and our method.
@@
Simple Ours Simple Ours Simple Ours Simple Ours
@@
An example of the spatial relationship task is presented in Figure 2-(d); To evaluate our method, we generated 2D scenes using the synthetic dataset described earlier and posed four-option multiple-choice questions for each scene, asking models to select the answer that best describes the spatial relation between the target objects. The evaluation results are shown in Table 4. As you can see, the new results show that our method (“Ours” column) yields consistent gains for most models: GPT-4o rises from 43.00% to 52.50% (+9.50%) on 2D, 64.00% to 68.50% (+4.50%) on 3D, and 69.39% to 77.43% (+8.04%) on natural scenes; Claude-sonnet jumps from 34.18% to 36.26% (+2.08%) on 2D, 72.50% to 76.50% (+4.00%) on 3D, and 37.43% to 46.15% (+8.72%) on natural images; LLaMa4 dips 2.00% on 2D but improves 3.50% on 3D and 8.17% on natural; and Qwen2.5-VL gains 1.50% on 2D and 4.5% on 3D yet sees a 3.06% drop on natural scenes. Overall, our approach particularly enhances 3D spatial reasoning—most dramatically for Claude-sonnet—and substantially boosts real-image performance in large models, while smaller or vision-augmented models show more mixed effects on natural data.
@@ 4.6 Chain of Thought
Figure 3: Comparison of our method with Chain-of-Thought (CoT) prompting on the GPT-4o model across four tasks. Each subplot shows the performance of three methods—Simple (base), CoT, and Ours (our method)—evaluated using a task-specific metric indicated on the x-axis. Bars are grouped into 2D and 3D datasets.
@@
To examine whether purely textual reasoning cues can bridge the binding gap, we pit our visually structured prompt against a Chain-of-Thought (CoT) baseline. The baseline employs GPT-4o with the canonical cue “Let’s think step by step” [13] appended to each instruction, leaving the image untouched, mirroring standard practice for eliciting step-wise reasoning in language models. Across all four tasks, our method consistently surpasses both the Simple model (no reasoning prompt) and the CoT variant in 2D and 3D scenarios (Figure 3). Crucially, the gap is far more pronounced on the 2D datasets. Flattened scenes crowd objects into a single plane, intensifying cross-object interference and making accurate binding harder. In these cluttered 2D settings, the lightweight spatial scaffold provides the model with an external, row-by-row traversal plan, producing markedly larger gains than CoT; in 3D scenes, the relative advantage narrows but persists. These results underline that verbal reasoning alone is not enough. A CoT prompt can encourage the model to explain its answer, but it cannot repair a noisy, globally pooled image embedding. Once early tokens are conditioned on entangled visual features, every subsequent “step” inherits the same misbindings. The scaffold, by contrast, prestructures the visual input so that each attended region contains fewer competing objects, allowing the language model to reason over cleaner, localized representations. Thus, reliable multimodal reasoning emerges only when textual chains of thought are paired with a minimal but crucial spatial guide.
4.7 Fine-Tuning for Visual Reasoning
@@
Figure 4: Comparison of our method with visual reasoning–finetuned model on the Qwen2.5-VL base model across four tasks. Each subplot shows the performance of three methods—Qwen-Simple (Qwen2.5-VL), Qwen-Ours (our method applied to Qwen2.5-VL), and Mulberry (Qwen2.5-VL finetuned for visual reasoning)—evaluated using a task-specific metric indicated on the x-axis Bars are grouped into 2D and 3D datasets.
To evaluate whether our lightweight method can match the benefits of model-level adaptation, we compare it to Mulberry [24], a Qwen2.5-VL variant fine-tuned for visual reasoning. As shown in Figure 4, we evaluate the base Qwen2.5-VL model, our method, and the fine-tuned Mulberry model across all tasks. Our method consistently improves over the base model and often matches or outperforms Mulberry, despite requiring no fine-tuning or access to model internals. For example, in 2D counting, Qwen-Ours reaches 41% accuracy, surpassing Mulberry’s 15%. These results show that addressing the binding problem through simple visual scaffolding and spatial prompting, without additional supervision or computational cost, can be as effective as training-based approaches for enhancing visual reasoning.

### v4

@@ 4.2 Counting
Dataset Objects Baseline VISER Baseline VISER Baseline VISER Baseline VISER
@@
An instance of this task is demonstrated in Figure 2-(c); To evaluate our method, we generated 2D and 3D images, using the synthetic data set described above, each containing between 10 and 20 instances of the target object to increase complexity. We measure performance using counting accuracy; the results are summarized in Table 1. As presented, our counting‐specific prompting yields substantial gains on synthetic benchmarks: in 2D scenes, GPT-4o jumps from 12.00% to 38.83%, Claude-sonnet from 8.50% to 10.67%, LLaMa4 from 6.33% to 19.10%, and Qwen2.5-VL surges from 5.83% to 40.83%; in 3D scenes, GPT-4o climbs from 15.00% to 31.00%, Claude-sonnet from 17.33% to 22.00%, LLaMa4 from 14.00% to 36.39%, and Qwen2.5-VL from 8.51% to 26.67%. On natural images, improvements are more modest—GPT-4o from 29.82% to 35.65%, Claude-sonnet from 2.09% to 6.42%, LLaMa4 from 29.22% to 31.65%—and Qwen2.5-VL sees a slight drop (18.91% to 17.29%). Together, these results underscore the effectiveness of task‐aware prompting in controlled settings while highlighting the remaining challenge of generalizing to real‐world imagery.
4.3 Visual Search
Table 2: Harmonic mean comparison for the Visual Search task, evaluating VISER against base models across GPT-4o, Claude3.5-sonnet, LLaMa4, and Qwen2.5-VL on 2D and 3D datasets with varying numbers of objects.
GPT-4o Claude-sonnet LLaMa4 Qwen2.5-VL
Dataset Objects Baseline VISER Baseline VISER Baseline VISER Baseline VISER
2D 20 0.71 0.91 0.55 0.82 0.00 0.06 0.31 0.43
30 0.50 0.80 0.54 0.70 0.00 0.00 0.28 0.37
40 0.25 0.64 0.18 0.59 0.00 0.00 0.14 0.23
50 0.46 0.55 0.10 0.55 0.00 0.00 0.47 0.56
Avg 0.48 0.73 0.34 0.66 0.00 0.02 0.30 0.40
3D 20 1.00 1.00 1.00 1.00 0.82 0.80 0.33 0.44
30 0.96 0.96 0.88 0.91 0.43 0.60 0.15 0.18
40 0.89 0.94 0.72 0.80 0.59 0.58 0.00 0.15
50 0.81 0.84 0.59 0.75 0.39 0.38 0.00 0.04
Avg 0.91 0.93 0.80 0.86 0.56 0.59 0.12 0.20
An instance of the visual search problem is presented in Figure 2-(a); we evaluate the proposed method using the synthetic dataset described earlier to generate 2D and 3D scenes containing between 20 and 50 objects of varying complexities. We assessed performance by measuring accuracy in correctly identifying the target’s presence or absence, distinguishing between “visible” searches (the object is explicitly in the scene) and “invisible” searches (the object is absent). This distinction is important because simpler LVLMs often default to predicting “false,” inflating their invisible‐object accuracy while failing to detect visible targets. After computing the accuracy for visible and invisible subtasks, we combined them via the harmonic mean to produce a final score.
$\text{Harmonic Mean}=\frac{2\cdot\text{Visible}\cdot\text{Invisible}}{\text{Visible}+\text{Invisible}}$ (1)
Table 2 compares our approach against baselines, showing VISER delivers consistent and substantial gains over the baseline across 2D and 3D scenes and for all evaluated LVLMs. In 2D settings, the average harmonic mean for GPT-4o climbs from 0.48 to 0.73, for Claude3.5-sonnet from 0.34 to 0.66, for LLaMa4 from 0.00 to 0.02, and for Qwen2.5-VL from 0.30 to 0.40, with the largest relative improvements occurring as scene complexity increases (e.g., at 40 objects GPT-4o rises from 0.25 to 0.64, and Claude3.5-sonnet rises from 0.18 to 0.59). In 3D scenes, VISER still yields gains to 0.93, and weaker models benefit notably (Claude3.5-sonnet: 0.80 → 0.86; LLaMa4: 0.56 → 0.59; Qwen2.5-VL: 0.12 → 0.20). These results confirm that our strategy not only corrects the “always-false” bias of simpler LVLMs in invisible-object tasks but also markedly enhances their ability to detect visible targets, achieving robust performance regardless of object count, visibility, or scene dimensionality.
@@ 4.4 Scene Description
The scene description task requires models to generate accurate textual narratives that describe the objects, their attributes, and their relationships within an image. This task challenges a LVLM’s ability to solve the binding problem. The model must not only detect features (e.g., “red”, “cube”, “blue”, “sphere”) but also correctly associate them with the corresponding objects in its textual output (e.g., “a red cube and a blue sphere”, not “a blue cube and a red sphere”). Errors, such as attribute swapping or incorrectly pairing features with objects, often arise in complex scenes. This typically occurs when multiple objects share some features but differ in others, which inherently challenges precise feature-object binding.
Table 3: Comparison of scene description performance (lower is better; measured as average distance) for VISER versus baseline across GPT-4o and Claude3.5-sonnet on 2D and 3D datasets with varying numbers of objects.
@@
Dataset Objects Baseline VISER Baseline VISER Baseline VISER Baseline VISER
@@
Figure 2-(b) shows an instance of the scene description task; we evaluate a synthetically generated dataset of 2D and 3D scenes, each containing between 10 and 20 objects to vary scene complexity. Model performance is quantified by the edit distance between the generated descriptions and the ground-truth annotations; we show more detail about this in Appendix B. This yields three outcome categories: (1) an exact match, (2) a single-attribute mismatch (either shape or color), and (3) a double-attribute mismatch (both shape and color). The full results are presented in Table 3. The greatest average gains (-0.81 in 2D and -1.06 in 3D) are driven precisely where the task’s difficulty peaks, showing that the technique not only endures extra clutter but thrives on it. This confirms that VISER is the most reliable option for reasoning with a high object count.
@@ 4.5 Spatial Relationship
In this task, the model is expected to identify and verify the relative positions of objects in a scene (e.g., “Is the red cube to the left of the blue sphere?”). This involves not only correctly binding intrinsic features (e.g., color, shape) to objects but also accurately associating each object with its spatial location, followed by a relational comparison. Failure to bind features properly can lead to misidentification of the objects being queried. Furthermore, if the LVLM cannot accurately associate each identified object with its precise location, its capacity to reason about their relative positions will be compromised. LVLMs often struggle with spatial reasoning due to their inability to maintain distinct, spatially grounded representations of multiple objects simultaneously [8].
Table 4: Accuracy (%) comparison of spatial relationship predictions among GPT-4o, Claude 3.5-Sonnet, LLaMA 4, and Qwen2.5-VL on 2D, 3D, and natural image datasets, using both baseline methods and VISER.
@@
Baseline VISER Baseline VISER Baseline VISER Baseline VISER
@@
An example of the spatial relationship task is presented in Figure 2-(d); To evaluate VISER, we generated 2D and 3D scenes using the synthetic dataset described earlier and posed four-option multiple-choice questions for each scene, asking models to select the answer that best describes the spatial relation between the target objects. The evaluation results are shown in Table 4. As you can see, the new results show that VISER yields consistent gains for most models: GPT-4o rises from 43.00% to 52.50% (+9.50%) on 2D, 64.00% to 68.50% (+4.50%) on 3D, and 69.39% to 77.43% (+8.04%) on natural scenes; Claude-sonnet jumps from 34.18% to 36.26% (+2.08%) on 2D, 72.50% to 76.50% (+4.00%) on 3D, and 37.43% to 46.15% (+8.72%) on natural images; LLaMa4 dips 2.00% on 2D but improves 3.50% on 3D and 8.17% on natural; and Qwen2.5-VL gains 1.50% on 2D and 4.5% on 3D yet sees a 3.06% drop on natural scenes.
@@ 4.6 Chain of Thought
Figure 3: Comparison of VISER with Chain-of-Thought (CoT) prompting on the GPT-4o model across four tasks. Each subplot shows the performance of three methods (Baseline, CoT, and VISER) evaluated using a task-specific metric indicated on the x-axis. Bars are grouped into 2D and 3D datasets.
@@
To examine whether purely textual reasoning cues can bridge the binding gap, we pit our visually structured prompt against a CoT baseline. The baseline employs GPT-4o with the canonical cue “Let’s think step by step” [13] appended to each instruction, leaving the image untouched, mirroring standard practice for eliciting step-wise reasoning in language models. Across all four tasks, VISER consistently surpasses both the Baseline model (no reasoning prompt) and the CoT variant in 2D and 3D scenarios (Figure 3). Crucially, the gap is far more pronounced on the 2D datasets. Flattened scenes crowd objects into a single plane, intensifying cross-object interference and making accurate binding harder. In these cluttered 2D settings, the lightweight spatial scaffold provides the model with an external, row-by-row traversal plan, producing markedly larger gains than CoT; in 3D scenes, the relative advantage narrows but persists. These results underline that verbal reasoning alone is not enough. A CoT prompt can encourage the model to explain its answer, but it cannot repair a noisy, globally pooled image embedding. Once early tokens are conditioned on entangled visual features, every subsequent “step” inherits the same misbindings. The scaffold, by contrast, prestructures the visual input so that each attended region contains fewer competing objects, allowing the language model to reason over cleaner, localized representations. Thus, reliable multimodal reasoning emerges only when textual chains of thought are paired with a minimal but crucial spatial guide.
4.7 Comparison with Fine-Tuned Models for Visual Reasoning
@@
Figure 4: Comparison of VISER with visual reasoning–finetuned model on the Qwen2.5-VL base model across four tasks. Each subplot shows the performance of four methods: Qwen-Baseline (Qwen2.5-VL), Qwen-VISER (VISER applied to Qwen2.5-VL), Mulberry (Qwen2.5-VL finetuned for visual reasoning), and OpenVLThinker (RL-finetuned Qwen2.5-VL), evaluated using task-specific metrics, and results are grouped into 2D and 3D datasets.
To evaluate whether our lightweight method can match the benefits of model-level adaptation, we compare it to Mulberry[24]—a Qwen2.5-VL variant fine-tuned for visual reasoning—and OpenVLThinker[43], which enhances reasoning in a Qwen backbone via reinforcement learning using a GRPO implementation. As illustrated in Figure 4, we evaluate the base Qwen2.5-VL model, VISER, the fine-tuned Mulberry, and OpenVLThinker. VISER consistently improves upon the base model and often matches or outperforms fine-tuned models, despite requiring neither fine-tuning nor access to model internals.
For example, on the 2D-Counting task, VISER achieves a significant boost in performance—41% accuracy, compared to 15% by Mulberry and the same result from OpenVLThinker. The only tasks where OpenVLThinker shows an advantage are 2D-Spatial reasoning and 3D-Counting, likely due to its reinforcement learning-based optimization for multimodal reasoning. This suggests that while training-intensive methods can excel in specific domains, they incur substantial computational costs.
In contrast, our training-free paradigm demonstrates broader effectiveness across diverse multimodal reasoning tasks, offering superior scalability and efficiency. These results show that addressing the binding problem through simple visual scaffolding and spatial prompting (without additional supervision or computational expense) can rival and even surpass state-of-the-art training-based techniques for multimodal reasoning.
4.8 Broader Evaluation Across Reasoning Benchmarks
To further validate the robustness and generality of our method, we evaluate it on additional reasoning-focused benchmarks, even in domains where visual structural cues might seem less relevant. These benchmarks include MMBench [44], PhysBench [45], RAVEN [46], and a Visual Analogy task [9]. These evaluations span a broad range of visual reasoning challenges, covering attribute, logical, relational, physical, and analogical reasoning. Results consistently show that our method improves performance across diverse tasks and modalities, demonstrating strong generalization in both abstract and grounded visual reasoning scenarios.
MMBench
MMbench is a reasoning-focused benchmark composed of tasks in Attribute, Logical, and Relation Reasoning. We include all reasoning tasks in our evaluation. As shown in Table 5, our method surpasses the baseline across most categories using both GPT-4o and Qwen2.5-VL models.
Table 5: Performance comparison (Accuracy %) of VISER vs. baseline on the MMBench benchmark. Results are reported across three key reasoning categories: Attribute, Logical, and Relation.
Model Method Attribute Reasoning Logical Reasoning Relation Reasoning
GPT-4o Baseline 88.67 78.25 75.17
VISER 88.00 82.75 81.12
Qwen2.5-VL Baseline 80.67 80.00 82.99
VISER 82.33 80.00 83.33
PhysBench
We also evaluate on PhysBench, a benchmark focused on physical commonsense reasoning with categories including Dynamics, Scene Understanding, Object Relationships, and Object Properties. Table 6 presents results on the validation split. Our approach consistently improves performance across most categories, showcasing its utility in physical reasoning tasks.
Table 6: Comparison of VISER and the baseline on the PhysBench benchmark, covering physical reasoning tasks across four categories: Dynamics, Scene Understanding, Object Relationships, and Object Property. Overall accuracy is also reported. All values are reported as Accuracy (%).
Model Method Dynamics Scene Understanding Object Relationships Object Property Overall
GPT-4o Baseline 35.00 33.33 75.00 55.00 55.68
VISER 40.00 41.67 75.00 70.00 61.37
Qwen2.5-VL Baseline 40.00 50.00 61.11 60.00 54.55
VISER 40.00 50.00 72.22 70.00 61.36
RAVEN
We evaluate our method on the RAVEN benchmark, which already includes grid lines. Instead of altering the images, we added the prompt: “Scan the image sequentially based on the grid lines present in the image” to encourage the model to leverage existing structure. Using GPT-4o, our method correctly answered 26 out of 140 questions, compared to 19 for the baseline—highlighting that prompting alone, without image alteration, can yield tangible gains when leveraging structural cues.
Visual Analogy Task
We evaluate the Visual Analogy task under a unified single-image setup. Using GPT-4o, both the baseline and our method achieve perfect accuracy (100%). For Qwen2.5-VL, our method improves performance from 72.0% to 77.0%, showing benefits in abstract analogical reasoning.